An 14.0-mu F capacitor is connected across an AC s

R3
Answer #1

given

C = 14 micro F

Vmax = 80 v

f = 65 hz

capacitive reactance, XC = 1/(2*pi*f*C)

= 1/(2*pi*65*14*10^-6)

= 174.9 ohms

so, Imax = Vmax/XC

= 80/174.9

= 0.457 A or 457 mA